OpenStudy (anonymous):

simplify the exponential expressions (6x^2)^3x^-15 (x^3)^3/x^15

jhonyy9 (jhonyy9):

use (3x^a)^b = (3^b)*x^(a*b) and x^(-a) = 1/(x^a) x^a /x^b = x^(a-b)